OpenStudy (anonymous):

Jesse played two days of golf. On the second day he got a score of -2. His total for the two days was -8. What was jesses score on the first day? Choose a label a variable for this situation. Write a equation using the variable that would answer the question. Solve your equation and check your answer. Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

dont post the same question over a number of times i remember i had answered this question with full explanation

OpenStudy (anonymous):

This is my first time posting this question. Must of been someone else.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Choose a label a variable for this situation. Write a equation using the variable that would answer the question. Solve your equation and check your answer. Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (anonymous):

Sheg wouldnt help me

OpenStudy (anonymous):

x+(-2)=8

OpenStudy (anonymous):

Now solve for x

OpenStudy (anonymous):

Whoops x-2=-8

OpenStudy (turingtest):

are you happy with rld's answer? it is correct, but does not go step-by-step as the problem requests

OpenStudy (anonymous):

x represents the first day's score and -2 is the second days score

OpenStudy (anonymous):

@Turningtest but she helps me and then she goes into steps.

OpenStudy (anonymous):

and both of them together =-8

OpenStudy (turingtest):

perfect :)

OpenStudy (anonymous):

Dont worry abt turingtest he is really nice

OpenStudy (anonymous):

Always helps me out step by steo

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Choose a label a variable for this situation. Write a equation using the variable that would answer the question. Solve your equation and check your answer. Use the solution to write a sentence answer to the question in the word problem.

OpenStudy (anonymous):

ok so we gave the equation

OpenStudy (anonymous):

and solve it what does x equal?

OpenStudy (anonymous):

Then plug in the answer you get into th eoriginal equation to see if it correct

OpenStudy (anonymous):

Jess's score on the first day was -2

OpenStudy (anonymous):

I think i covered e/t. did u follow?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok but remember to plug -2 into the original equation to show that the answer you got is correct

OpenStudy (anonymous):

I did :) What does X equal?

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

remeber how we solved it

OpenStudy (anonymous):

whoops x=-6

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

x-2+2=-8+2 x=-6

OpenStudy (anonymous):

so plug in -6 into the original equation again to show that it is correct